OpenStudy (anonymous):

@mpj4 7. which method would be the simplest way to solve the system? y=1/2x 2x+3y=28 a. graphing b. substitution c. elimination d. distributive 8. what is the solution of the system? y=-3x 3x+2y=6 a. (2,6) b. (-1,3) c. (-2,6) d. (1.5, -4.5) 9. For the following system of equations, what is the x-value of the solution? -x+2y=6 6y=x+18 a. -1 b. 0 c. 1 d. 6

OpenStudy (igreen):

You could substitute 1/2x for y in the 2nd equation.

OpenStudy (igreen):

Graphing would be the easiest to me, but I'm betting they want you to go for substitution because it already gives you what y is.

OpenStudy (mpj4):

7 B 8 C 9 D

OpenStudy (igreen):

@mpj4 Don't give answers.

OpenStudy (mpj4):

Ahh, so that's how it works.

OpenStudy (igreen):

For the 2nd one, we can also substitute -3x for y into the 2nd equation. \(3x + 2y = 6\) \(3x + 2(-3x) = 6\) Multiply: \(3x - 6x = 6\) Subtract: \(-3x = 6\) Divide -3 to both sides. Can you divide 6 / -3? @Bpgrace